# On finding optimal collective variables for complex systems by minimizing the deviation between effective and full dynamics

### v1

@@ Appendix A Transfer operator for Langevin dynamics
Appendix B Rigorous definition of $\mu_{z}$
In this section, we provide rigorous definition of the conditional probability measure $\mu_{z}$. Given the CV map $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ and $z\in\mathbb{R}^{k}$, recall the level set $\Sigma_{z}:=\{x\in\mathbb{R}^{d},\xi(x)=z\}$, which is a $(d-k)$-dimensional submanifold of $\mathbb{R}^{d}$ under certain assumptions on $\xi$. Given the probability measure $\mu$ in $\mathbb{R}^{d}$ with probability density $\pi$, the probability measure $\mu_{z}$ in (39) can be rigorously defined as a probability measure on $\Sigma_{z}$ by
$\mu_{z}(\mathrm{d}{x})=p_{z}(x)\sigma_{\Sigma_{z}}(\mathrm{d}{x})=\frac{1}{Q(z)}\pi(x)[\mathrm{det}(\nabla\xi(x)\nabla\xi(x)^{\top})]^{-\frac{1}{2}}\,\sigma_{\Sigma_{z}}(\mathrm{d}{x})\,,$ (85)
where $\sigma_{\Sigma_{z}}$ denotes the surface measure of $\Sigma_{z}$ and
$Q(z)=\int_{\Sigma_{z}}\pi(x)[\mathrm{det}(\nabla\xi(x)\nabla\xi(x)^{\top})]^{-\frac{1}{2}}\,\sigma_{\Sigma_{z}}(\mathrm{d}{x})$ (86)
is the rigorous definition of the normalizing constant in (40). A determinant term is introduced in (85) and (86), so that by co-area formula [12, Theorem 3.11] we have, for all integrable functions $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$,
$\int_{\mathbb{R}^{d}}f(x)\pi(x)\,\mathrm{d}{x}=\int_{\mathbb{R}^{k}}\Big{(}\int_{\Sigma_{z}}f(x)\,\mu_{z}(\mathrm{d}{x})\Big{)}Q(z)\,\mathrm{d}{z}\,.$ (87)
The measure $\mu_{z}$ is called the conditional probability measure, since it holds
for all integrable functions $f$.

### v2

@@ Appendix A Transfer operator for Langevin dynamics
Appendix B Proofs in Section 2
In this section, we present the proofs of Proposition 1 and Theorem 2.2 in Section 2.2, as well as the proof of Proposition 3 in Section 2.3.
Proof of Proposition 1.
To prove the first claim, let us apply [51, Theorem VI.23] and show that $\mathcal{T}$ is a Hilbert-Schmidt operator if (21) holds. Obviously, by the definition (2) we have
Therefore, the conditions of [51, Theorem VI.23] hold and it follows that $\mathcal{T}$ is a Hilbert-Schmidt operator. This in turn implies that $\mathcal{T}$ is compact and any nonzero element in its spectrum $\sigma(\mathcal{T})$ is an eigenvalue (see [51, Theorem VI.22] and [51, Theorem VI.15]).
Next, we prove the second claim. Assume without loss of generality that the eigenfunction $\varphi$ corresponding to the eigenvalue $\lambda$ is normalized such that $\langle\varphi,\varphi\rangle_{\mu}=1$. Multiplying both sides of (20) by $\varphi$ and integrating with respect to $\mu$, we get $\lambda=\langle\mathcal{T}\varphi,\varphi\rangle_{\mu}$. Concerning the upper bound in the second claim, using (11)–(12) we can derive that $0\leq\mathcal{E}(\varphi)=\langle(I-\mathcal{T})\varphi,\varphi\rangle_{\mu}=1-\lambda$, which implies $\lambda\leq 1$. If $\lambda=1$, we must have $\mathcal{E}(\varphi)=0$ and therefore (11) implies that $\varphi$ is constant, since both $p$ and $\pi$ are positive. For the lower bound, similar to the proof of (12) in Lemma 13, we have (since $\mathcal{T}$ is self-adjoint)
$0\leq\frac{1}{2}\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d}}(\varphi(y)+\varphi(x))^{2}p(x,y)\mathrm{d}{y}\right)\,\mu(\mathrm{d}{x})=\langle(I+\mathcal{T})\varphi,\varphi\rangle_{\mu}=1+\lambda\,,$
Proof of Theorem 2.2.
Let $f_{1},f_{2},\dots,f_{m}\in\mathcal{H}$ be $m$ functions such that (26) holds. Define the diagonal matrix $A=\mbox{diag}\{\omega_{1},\omega_{2},\dots,\omega_{m}\}$ and the $m\times m$ symmetric matrix $B=(B_{ij})_{1\leq i,j\leq m}$, whose entries are $B_{ij}=\mathcal{E}(f_{i},f_{j})$ (see (10)). Using the fact that $A$ is diagonal and $B_{ii}=\mathcal{E}(f_{i},f_{i})=\mathcal{E}(f_{i})$ for $1\leq i\leq m$ (see (11)), we have
$\sum_{i=1}^{m}\omega_{i}\mathcal{E}(f_{i})=\mbox{{tr}}(AB)\,.$ (90)
Also, we have
$c^{\top}Bc=\sum_{1\leq i,j\leq m}c_{i}B_{ij}c_{j}=\sum_{1\leq i,j\leq m}c_{i}\mathcal{E}(f_{i},f_{j})c_{j}=\mathcal{E}\Big{(}\sum_{i=1}^{m}c_{i}f_{i}\Big{)}\geq 0\,,\quad\forall c\in\mathbb{R}^{m}\,,$ (91)
and it equals zero only if $c=0$ (because (26) implies that $f_{1},f_{2},\cdots,f_{m}$ are linearly independent and $\mathcal{E}$ is defined in (11) with positive $p$). This implies that $B$ is a positive definite symmetric matrix, whose eigenvalues are denoted by $0<\widetilde{\lambda}_{1}\leq\widetilde{\lambda}_{2}\leq\dots\leq\widetilde{\lambda}_{m}$. Applying Ruhe’s trace inequality [41, H.1.h, Section H, Chapter 9] and using the assumption that $\omega_{1}\geq\dots\geq\omega_{m}>0$, we obtain from (90) that
$\sum_{i=1}^{m}\omega_{i}\mathcal{E}(f_{i})=\mbox{{tr}}(AB)\geq\sum_{i=1}^{m}\omega_{i}\widetilde{\lambda}_{i}\,.$ (92)
Let us show that $\widetilde{\lambda}_{j}\geq 1-\lambda_{j}$ for $j\in\{1,2,\dots,m\}$. For this purpose, applying the min-max theorem for symmetric matrices we have
$\widetilde{\lambda}_{j}=\min_{S_{j}}\max_{c\in S_{j},|c|=1}c^{\top}Bc=\min_{S_{j}}\max_{c\in S_{j},|c|=1}\mathcal{E}\big{(}\sum_{i=1}^{m}c_{i}f_{i}\big{)}\,,$ (93)
where $S_{j}$ goes over all $j$-dimensional linear subspaces of $\mathbb{R}^{m}$ and the second equality follows from (91). Since $(f_{i})_{1\leq i\leq m}\subset\mathcal{H}$ satisfy (26), the functions $\mathbf{1},f_{1},\dots,f_{m}$ are pairwise orthonormal, where $\mathbf{1}$ denotes the constant function that equals one. Hence, each $j$-dimensional linear subspace $S_{j}\subset\mathbb{R}^{m}$ defines a $(j+1)$-dimensional linear subspace of $\mathcal{H}$ by $H_{j}=\big{\{}c_{0}\mathbf{1}+\sum_{i=1}^{m}c_{i}f_{i}\,|\,c_{0}\in\mathbb{R},c\in S_{j}\big{\}}$, such that $H_{j}\subset\mbox{span}\{\mathbf{1},f_{1},f_{2},\dots,f_{m}\}$. For $f=c_{0}\mathbf{1}+\sum_{i=1}^{m}c_{i}f_{i}$, the pairwise orthonormality and the fact $\mathcal{E}(f)=\mathcal{E}(f-c_{0})$ (see (11)) imply that $\|f\|_{\mu}=c_{0}^{2}+|c|^{2}$ and $\mathcal{E}(f)=\mathcal{E}\big{(}\sum_{i=1}^{m}c_{i}f_{i}\big{)}$. Hence, we have
$\max_{f\in H_{j},\|f\|_{\mu}=1}\mathcal{E}(f)=$ $\max_{\begin{subarray}{c}c_{0}\in\mathbb{R},c\in S_{j},c_{0}^{2}+|c|^{2}=1\end{subarray}}\mathcal{E}(\sum_{i=1}^{m}c_{i}f_{i})$
$=$ $\max_{\begin{subarray}{c}c\in S_{j},|c|\leq 1\end{subarray}}\mathcal{E}(\sum_{i=1}^{m}c_{i}f_{i})=\max_{\begin{subarray}{c}c\in S_{j},|c|=1\end{subarray}}\mathcal{E}(\sum_{i=1}^{m}c_{i}f_{i})\,,$
where the last equality follows from the fact that $\mathcal{E}$ is a (nonnegative) quadratic form. Therefore, from (93) we find that
$\begin{split}\widetilde{\lambda}_{j}=&\min_{S_{j}}\max_{c\in S_{j},|c|=1}\mathcal{E}\big{(}\sum_{i=1}^{m}c_{i}f_{i}\big{)}\\ =&\min_{H_{j}}\max_{f\in H_{j},\|f\|_{\mu}=1}\mathcal{E}(f)\\ =&\min_{H_{j}}\max_{f\in H_{j},\|f\|_{\mu}=1}\langle(\mathcal{I}-\mathcal{T})f,f\rangle_{\mu}\\ \geq&1-\lambda_{j}\,,\end{split}$ (94)
where the last inequality follows from the min-max Theorem [56, Theorem 4.14] and the fact that $1-\lambda_{j}$ is the $(j+1)$th smallest eigenvalue of $\mathcal{I}-\mathcal{T}$.
Combining (94) and (92), gives $\sum_{i=1}^{m}\omega_{i}\mathcal{E}(f_{i})\geq\sum_{i=1}^{m}\omega_{i}(1-\lambda_{i})$. Since the eigenfunctions $(\varphi_{i})_{1\leq i\leq m}$ satisfy (26) and $\sum_{i=1}^{m}\omega_{i}\mathcal{E}(\varphi_{i})=\sum_{i=1}^{m}\omega_{i}(1-\lambda_{i})$, we conclude that the identity (25) holds and the minimum is achieved when $f_{i}=\varphi_{i}$ for $i\in\{1,2,\dots,m\}$. ∎
Proof of Proposition 3.
$M_{N}^{R}=\sum_{m=0}^{N-1}\mathbbm{1}_{A}(X_{m})\mathbbm{1}_{B}(X_{m+j_{m}})\,,$
$k_{AB}=\lim_{N\rightarrow+\infty}\frac{1}{N}\sum_{m=0}^{N-1}\mathbbm{1}_{A}(X_{m})\mathbbm{1}_{B}(X_{m+j_{m}})=\int_{A}\left(\int_{\mathbb{R}^{d}}p(x,y)q(y)\,\mathrm{d}{y}\right)\mu(\mathrm{d}{x})\,,$
where the second equality follows from the definition of the committor $q$ in (31), the ergodicity of the process $X_{n}$ and Birkhoff’s ergodic theorem (see [31, Theorem 4.4 in Chapter 1]). Similarly, by considering reactive segments from $B$ to $A$ and applying the identity in the first claim, we obtain the second equality in this claim, since
$=$ $\int_{B}\left(\int_{\mathbb{R}^{d}}p(x,y)(1-q(y))\,\mathrm{d}{y}\right)\mu(\mathrm{d}{x})\,.$
Concerning the third claim, from (11) we can compute
$\mathcal{E}(q)=$ $\frac{1}{2}\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d}}|q(y)-q(x)|^{2}p(x,y)\mathrm{d}{y}\,\right)\mu(\mathrm{d}{x})$
$=$ $\langle q,(\mathcal{I}-\mathcal{T})q\rangle_{\mu}$
$=$ $\int_{A}q(x)[(\mathcal{I}-\mathcal{T})q](x)\mu(\mathrm{d}{x})+\int_{(A\cup B)^{c}}q(x)[(\mathcal{I}-\mathcal{T})q](x)\mu(\mathrm{d}{x})$
$+\int_{B}q(x)[(\mathcal{I}-\mathcal{T})q](x)\mu(\mathrm{d}{x})$
$=$ $\int_{B}[(\mathcal{I}-\mathcal{T})q](x)\mu(\mathrm{d}{x})$
$=$ $\int_{B}\left(\int_{\mathbb{R}^{d}}p(x,y)(1-q(y))\,\mathrm{d}{y}\right)\mu(\mathrm{d}{x})$
$=$ $k_{AB}\,,$
where the second equality follows from (12) in Lemma 13 and the fact that $\langle f,(\mathcal{I}-\mathcal{T}^{rev})f\rangle_{\mu}=\langle f,(\mathcal{I}-\mathcal{T})f\rangle_{\mu}$ for any $f$, the fourth equality follows by using the equation (32) of the committor $q$, the definition of $\mathcal{T}$ and the boundary conditions of $q$ on $A,B$, the fifth equality follows from the fact that $\int_{\mathbb{R}^{d}}p(x,y)\mathrm{d}{y}=1$ and $q|_{B}=1$, and the last equality follows from (95). ∎